# GHZ protocols enhance frequency metrology despite spontaneous decay

### v1

@@ Abstract
The use of correlated states and measurements promises improvements in the accuracy of frequency metrology and the stability of atomic clocks. However, developing strategies robust against dominant noise processes remains challenging. We address the issue of decoherence due to spontaneous decay and show that Greenberger-Horne-Zeilinger (GHZ) states, in conjunction with a correlated measurement and nonlinear estimation strategy, achieve gains comparable to fundamental bounds for ensembles of up to 40 atoms. This result is surprising since GHZ states do not provide any enhancement under dephasing noise compared to the standard quantum limit of uncorrelated states. The gain arises from a veto signal, which allows for the detection and mitigation of errors caused by spontaneous emission events. Through comprehensive Monte-Carlo simulations of atomic clocks, we demonstrate the robustness of the GHZ protocol.
@@
Unfortunately, decoherence presents a significant obstacle in frequency metrology, impairing the precision of measurements by compromising the coherence of quantum systems essential for achieving entanglement-based enhancement [40, 41, 42, 43, 44, 45]. Among technical noise sources, residual fluctuations of the laser and magnetic field noise impose major limitations [1, 2, 3, 46, 4, 47]. In particular, it has been demonstrated that GHZ protocols, which are optimal in the absence of decoherence, are highly sensitive to dephasing due to frequency fluctuations and ultimately show no improvement over SQL [44]. To address this susceptibility of entanglement-enhanced protocols, frequency fluctuations were taken into account to determine optimal interrogation sequences [48, 49, 50] as well as multi-ensemble schemes to effectively extend the coherence time were proposed recently [51, 52, 39, 53, 54, 55, 56, 57].
@@
Figure 1: Optimal GHZ protocol embedded in the full feedback loop of an atomic clock. A local oscillator (LO) with fluctuating frequency $\omega_{\mathrm{LO}}$ is stabilized to an atomic transition $\omega_{0}$. During the free evolution time $T$, the probe state accumulates a phase $\phi$ and is subject to spontaneous emission with rate $\Gamma$. Based on the measurement outcome $x$ of the observable $X$, an estimation $\phi_{\mathrm{est}}$ of the phase is conducted to corrected the LO accordingly.
@@
In contrast to the extensive treatment of frequency fluctuations, the effects of spontaneous emission have received comparatively little attention. However, the finite lifetime of the qubits in the excited state represents a fundamental limit rather than an external noise source. With state-of-the-art clock lasers reaching coherence times of several seconds [58], the excited-state lifetime of various clock candidates becomes the more stringend limitation, for instance Ca+-ions ($1.1\,\mathrm{s}$), Sr+-ions ($0.4\,\mathrm{s}$), In+-ions ($0.2\,\mathrm{s}$) and Hg-atoms ($1.6\,\mathrm{s}$). With further technological improvements in the short-term stabilization of lasers, even clock species as Al+-ions ($20.7\,\mathrm{s}$) and Yb-atoms ($15.9\,\mathrm{s}$) will enter this regime [4]. This applies especially when using strongly entangled states which are generally more susceptible to decoherence e.g. due to spontaneous emission. Consequently, it is highly relevant to investigate the impact of spontaneous decay for the development of future clocks and identify optimal interrogation schemes for particular setups.
@@
Here, we will present a protocol with quantum operations of low complexity and a highly nonlinear estimator that saturates the quantum Cramér-Rao bound (QCRB) of the GHZ state (cf. Fig. 1). Surprisingly, and in contrast to dephasing, we find that GHZ states provide a significant enhancement compared to the SQL in the presence of spontaneous decay. Moreover, we will compare the sensitivity of this protocol to the ultimate lower limit and squeezed spin states (SSS) which are optimal in the asymptotic limit. Finally, we will show the robustness of the measurement and estimation scheme by pe rforming Monte-Carlo simulations of the full feedback loop in an atomic clock. Note that although we focus on the specific application to atomic clocks, the underlying concept extends to Ramsey interferometry in general, as utilized in various precision measurements, such as atom interferometry and magnetometry.
@@
where we fix the total averaging time $\tau=nT$ and denote $\Delta\phi(T)$ as the phase estimation uncertainty of a single Ramsey interrogation (cf. Methods). Disregarding decoherence, $\Delta\phi$ is independent of time, and the frequency estimation uncertainty appears to monotonically decrease for longer interrogation times $T$. If $N$ atoms are prepared in uncorrelated states, the phase uncertainty is limited by the standard quantum limit (SQL) $\Delta\phi^{2}\geq\Delta\phi_{\mathrm{SQL}}^{2}=1/N$ [2], such that $\Delta\omega(T)^{2}\geq(\Delta\omega_{\mathrm{SQL}}(T))^{2}=1/\tau TN$. This bound can be saturated with atoms prepared in a coherent spin state (CSS) and measurement of the spin projection along a suitable direction. A dramatic improvement can, in principle, be gained by preparing atoms in a GHZ state $|\mathrm{GHZ}\rangle=\left(|\negmedspace\uparrow\rangle^{\otimes N}+|\negmedspace\downarrow\rangle^{\otimes N}\right)/\sqrt{2}$ and measuring the parity $\Pi=(-1)^{N}\sigma_{x}^{\otimes N}$ after phase imprint during a Ramsey interrogation, as initially proposed by Wineland et al. [72]. This strategy, to which we refer in the following as ’parity-GHZ’ protocol, surpasses the SQL and saturates the Heisenberg limit (HL) $(\Delta\phi_{\mathrm{HL}})^{2}=1/N^{2}$ [2]. The improvement is due to the signal in Ramsey interference oscillating $N$ times faster, thus enhancing the sensitivity by exactly this factor.
@@
Indeed, performing a parity measurement on the GHZ state in the presence of spontaneous decay is not optimal, i.e. does not saturate the quantum Cramér-Rao bound (QCRB). We show in the Supplementary Materials that the QCRB for the phase uncertainty is given by
@@
The corresponding bound for the optimal frequency uncertainty follows from minimizing with respect to the interrogation time, $(\Delta\omega_{\mathrm{QCRB}}^{\mathrm{GHZ}})^{2}=\min_{T}(\Delta\phi_{\mathrm{QCRB}}^{\mathrm{GHZ}}(T))^{2}/T\tau$. For vanishing spontaneous emission, $\Gamma=0$, this reproduces the result of Huelga et al. [44] that the QCRB equals the SQL. However, for $\Gamma>0$ one finds $(\Delta\omega_{\mathrm{QCRB}}^{\mathrm{GHZ}})^{2}<(\Delta\omega_{\mathrm{SQL}})^{2}$ which is illustrated in Fig. 2(a). Somewhat surprisingly, GHZ states do admit gains beyond the SQL when the relevant decoherence process is spontaneous emission rather than dephasing noise. Since these gains are not realized by parity measurements, the question arises as to which alternative measurements do saturate the QCRB of the GHZ state (3).
@@
$\mathcal{U}_{\mathrm{GHZ}}$ =$=\begin{cases}\mathcal{T}_{x}(\pi)&\text{if }N\text{ is even}\\ \mathcal{R}_{x}\left(\frac{\pi}{2}\right)\mathcal{T}_{x}(\pi)&\text{if }N\text{ is odd}\end{cases}$ (4)
where $\mathcal{T}_{x}(\mu)=\exp\left(-i\frac{\mu}{2}S_{x}^{2}\right)$ denotes an OAT operation with twisting strength $\mu=\pi$ along $x$, whereas $\mathcal{R}_{x}(\theta)$ represents a rotation around the $x$-axis by an angle $\theta=\frac{\pi}{2}$. We note that the unitary $\mathcal{U}_{\mathrm{GHZ}}$ also corresponds to the operation that may be used to generate the GHZ state from the ground state $|\negmedspace\!\downarrow\rangle^{\otimes N}$ initially. Subsequently, atoms are measured projectively along $z$, effectively measuring the operator $X=\mathcal{U}_{\mathrm{GHZ}}\,S_{z}\,\mathcal{U}_{\mathrm{GHZ}}^{\dagger}$ with outcomes $x\in\{-\frac{N}{2},\ldots,\frac{N}{2}\}$. This procedure essentially corresponds to the one explored in the experiment reported by Leibfried et al. [66]. Any such measurement has to be complemented by a suitable rule for estimating the phase $\phi$ for a given outcome $x$. We can show that the nonlinear estimator
$\phi_{\mathrm{est}}(x,T)=$±$\begin{cases}\pm\mathrm{e}^{\frac{(\Gamma+\gamma)}{2}NT}&\text{if}x=\pm\frac{N}{2}\\ 0&\text{else}\end{cases}$ (5)
@@
The fact that this measurement and estimation strategy performs well under spontaneous emission can be understood from the two following reasons: Firstly, measurement outcomes $x\neq\pm\frac{N}{2}$ can occur only if at least one spontaneous emission event happened, such that selecting the $x=\pm\frac{N}{2}$ events, as is done by the estimator (5), in essence filters out all such cases. Secondly, and quite remarkably, for the particular measurement $X$ introduced above, the conditional probabilities $P(x|\phi,T)$ turn out to be independent of $\phi$ for all $x\neq\pm\frac{N}{2}$ (cf. Methods). Thus, these cases don’t deliver any information on $\phi$. Together, these two features allow to implement an error detection and mitigation scheme [73] designated to frequency metrology. On the one hand, outcomes $x\neq\pm\frac{N}{2}$ can be considered as heralded errors signifying an unsuccessful, decohered Ramsey interrogation. Conversely, outcomes $x=\pm\frac{N}{2}$ signal a no-jump dynamics delivering maximal phase information. On the other hand, exclusively selecting these events during (classical) post-processing results in enhanced sensitivity. We therefore refer to this scheme as ’heralded-GHZ’ protocol in the following.
@@
Figure 2: (a) Generic frequency estimation uncertainty $\Delta\omega(T)$ in the presence of spontaneous decay for $N=8$ scaled to the SQL (2). Symbols indicate the optimal interrogation time $T_{\mathrm{min}}$. (b) Relative frequency estimation uncertainty at the optimal interrogation time $T_{\mathrm{min}}$ for the indicated protocol types (see main text). The ratio $\Delta\omega/\Delta\omega_{\mathrm{SQL}}$ is independent of the decay rate $\Gamma$ due to the comparison with the SQL (triangle in (a)). Square and diamond symbols correspond to the minima in (a).
@@
Optimization over arbitrary entangled initial states, measurements, and estimators determines the performance of an optimal quantum interferometer (OQI). In the absence of decoherence, the OQI achieves the Heisenberg limit of sensitivity. Unfortunately, in the presence of decoherence, no general expressions for the OQI sensitivity for arbitrary ensemble sizes are available, but they rather require complex optimization procedures. An efficient algorithm is presented in Ref. [74], which allows for the computation of the OQI by iteratively optimizing the initial probe state and the observable, at least for small and moderate number of atoms $N$. However, numerical optimization becomes challenging with increasing ensemble size $N$. In the asymptotic limit ($N\gg 1$), lower bounds can be derived for several noise processes as discussed in Refs. [43, 42, 44, 75, 76, 77]. The ultimate lower bound in the presence of dephasing [44, 43, 42] and spontaneous decay [77] is given by
@@
which yields a maximal improvement of $1/\mathrm{e}$ over the SQL (2) and is asymptotically saturated by SSS (cf. Supplementary Materials). As a consequence, we will focus on small ensembles where the OQI can be evaluated numerically and benchmark the heralded-GHZ protocol against SSS generated by OAT since they are optimal in the asymptotic limit. In the following, we will consider setups where spontaneous decay is the main limiting decoherence process, and neglect dephasing noise assuming $\gamma=0$.
@@
In Fig. 2(b), we present the sensitivity of frequency estimation for a given particle number $N$, optimized in each case with respect to the interrogation time $T$ and scaled to the SQL. The conclusions we draw from Fig. 2(b) are: (i) The heralded-GHZ protocol shows a significant enhancement over the SQL for all $N$. In particular, a constant gain with regard to the the SQL is achieved for $N\geq 4$ and thus, no loss in improvement is observed for large $N$. Especially considering the scaling of small and intermediate ensembles, the resulting increase in sensitivity is quite remarkable. (ii) The importance of the nonlinear estimator (5) becomes evident since otherwise (denoted as ’linear-GHZ’ protocol), i.e. using the standard linear estimator, the maximal gain is obtained for $N=3$, whereas for larger ensembles the sensitivity converges to the SQL (cf. Methods). (iii) Although the heralded-GHZ protocol provides a substantial enhancement compared to the SQL, it does not saturate the OQI, which likewise represents a constant improvement with respect to the SQL in the regime ($N\leq 30$) where the numerical evaluation is feasible. Furthermore, in this regime the heralded-GHZ protocol achieves sensitivities relatively close to the OQI, despite its low complexity. (iv) For larger ensembles ($N\geq 42$), SSS achieve a higher sensitivity than the heralded-GHZ protocol and ultimately approximate the lower bound (6) asymptotically. Consequently, the gain of the OQI over the SQL will increase likewise for larger ensembles. In summary, this implies that GHZ states can be used to overcome the SQL in the presence of spontaneous emission and in addition perform close to the OQI for small and moderate ensemble sizes.
Figure 3: Monte-Carlo simulations of the full feedback loop in an atomic clock are displayed by symbols while theoretical predictions are shown as lines. Data points denote the average over $10$ independent clock runs. Error bars indicate the corresponding standard deviations. In each clock run, $10^{7}$ cycles were performed and the corresponding Allan deviation evaluated. The presented values are obtained by extrapolating the Allan deviation at $\tau=1~{}\mathrm{s}$ based on the long term stability (7) at $\tau\gg 1$.
@@
and thus corresponds to the frequency estimation uncertainty rescaled with $\omega_{0}$. Consequently, we present our results in two ways: Lower $x$-axis and left $y$-axis refer to general frequency estimation where the uncertainty is rescaled to be independent of the particular averaging time $\tau$ and lifetime $t_{\mathrm{spont}}=1/\Gamma$. Thus, it allows for an application to several experimental setups and atomic species as long as spontaneous decay remains the dominating decoherence effect. Upper $x$-axis and right $y$-axis illustrate results in an atomic clock for the particular example of Ca+-ions with lifetime $t_{\mathrm{spont}}=\frac{1}{\Gamma}\simeq 1.1\,\mathrm{s}$ [79] and transition frequency $\omega_{0}=2\pi\nu_{0}\simeq 411.042\,\mathrm{THz}$ [15]. Furthermore, we consider frequency fluctuations corresponding to a state of the art clock laser. In particular, we assume a laser limited by flicker noise with coherence time $Z\simeq 7.5~{}\mathrm{s}\gg t_{\mathrm{spont}}$ [58]. Results of numerical simulations (symbols) in comparison to theoretical predictions (lines) of the investigated protocols are shown in Fig. 3 for $N=4$ (top) and $N=16$ (bottom) particles. As Monte Carlo simulations are stochastic processes, resulting stabilities fluctuate around the average value. Overall, numerical simulations of all discussed interrogation schemes show very good agreement with theoretical predictions. Therefore, all schemes, including the heralded-GHZ protocol in particular, are robust and thus suited for realistic scenarios as in the context of atomic clocks. At long interrogation times $T\gtrsim t_{\mathrm{spont}}$, fringe hops occur where the feedback loop passes to an adjacent fringe resulting in the clock running systematically wrong and consequently spoiling the clock stability. Moreover, the comparison of $N=4$ and $N=16$ already indicates the transition between the optimality of GHZ-like protocols and squeezing protocols.
Conclusion & Outlook

### v2

@@ Abstract
The use of correlated states and measurements promises improvements in the accuracy of frequency metrology and the stability of atomic clocks. However, developing strategies robust against dominant noise processes remains challenging. We address the issue of decoherence due to spontaneous decay and show that Greenberger-Horne-Zeilinger (GHZ) states, in conjunction with a correlated measurement and nonlinear estimation strategy, achieve gains of up to 2.25 dB, comparable to fundamental bounds for up to about 80 atoms in the presence of decoherence. This result is surprising since GHZ states do not provide any enhancement under dephasing due to white frequency noise compared to the standard quantum limit of uncorrelated states. The gain arises from a veto signal, which allows for the detection and mitigation of errors caused by spontaneous emission events. Through comprehensive Monte-Carlo simulations of atomic clocks, we demonstrate the robustness of the GHZ protocol.
@@
Unfortunately, decoherence presents a substantial obstacle in frequency metrology, impairing the precision of measurements by compromising the coherence of quantum systems essential for achieving entanglement-based enhancement [39, 40, 41, 42, 43, 44]. Among technical noise sources, residual fluctuations of the laser and magnetic field noise impose major limitations [1, 2, 3, 45, 4, 46]. In particular, it has been demonstrated that GHZ protocols, which are optimal in the absence of decoherence, are highly sensitive to dephasing due to frequency fluctuations and ultimately show no improvement over SQL [43]. To address this susceptibility of entanglement-enhanced protocols, frequency fluctuations were taken into account to determine optimal interrogation sequences [47, 48, 49] as well as multi-ensemble schemes to effectively extend the coherence time were proposed recently [50, 51, 38, 52, 53, 54, 55, 56].
@@
Figure 1: Optimal GHZ protocol embedded in the full feedback loop of an atomic clock. A local oscillator (LO) with fluctuating frequency $\omega_{\mathrm{LO}}$ is stabilized to an atomic transition $\omega_{0}$. During the free evolution time $T$, the probe state accumulates a phase $\phi$ and is subject to spontaneous emission with rate $\Gamma$. Based on the measurement outcome $x$ of the observable $X$, an estimation $\phi_{\mathrm{est}}$ of the phase is conducted to correct the LO accordingly.
@@
In contrast to the extensive treatment of frequency fluctuations, the effects of spontaneous emission have received comparatively little attention. However, the finite lifetime of the qubits in the excited state represents a fundamental limit rather than an external noise source. With state-of-the-art clock lasers reaching coherence times of several seconds [57], the excited-state lifetime of various clock candidates becomes the more stringent limitation, for instance Ca+-ions ($1.1\,\mathrm{s}$), Sr+-ions ($0.4\,\mathrm{s}$), In+-ions ($0.2\,\mathrm{s}$) and Hg-atoms ($1.6\,\mathrm{s}$). With further technological improvements in the short-term stabilization of lasers, even clock species as Al+-ions ($20.7\,\mathrm{s}$) and Yb-atoms ($15.9\,\mathrm{s}$) will enter this regime [4]. This applies especially when using strongly entangled states which are generally more susceptible to decoherence e.g. due to spontaneous emission. Consequently, it is highly relevant to investigate the impact of spontaneous decay for the development of future clocks and identify optimal interrogation schemes for particular setups.
@@
Here, we will present a protocol with quantum operations of low complexity and a highly nonlinear estimator that saturates the quantum Cramér-Rao bound (QCRB) of the GHZ state (cf. Fig. 1). Surprisingly, and in contrast to dephasing, we find that GHZ states provide a substantial enhancement compared to the SQL in the presence of spontaneous decay. Moreover, we will compare the sensitivity of this protocol to the ultimate lower limit [58] and to squeezed spin states (SSS) which are optimal in the asymptotic limit for large particle number. In addition, we will present a variation of this protocol with a GHZ-like initial state reaching the level of the ultimate lower limit for up to about 80 atoms. Finally, we will show the robustness of the measurement and estimation scheme by performing Monte-Carlo simulations of the full feedback loop in an atomic clock. Note that although we focus on the specific application to atomic clocks, the underlying concept extends to Ramsey interferometry in general, as utilized in various precision measurements, such as atom interferometry and magnetometry.
@@
where we fix the total averaging time $\tau=nT$ and denote $\Delta\phi(T)$ as the phase estimation uncertainty of a single Ramsey interrogation (cf. Methods). Disregarding decoherence, $\Delta\phi$ is independent of time, and the frequency estimation uncertainty appears to monotonically decrease for longer interrogation times $T$. If $N$ atoms are prepared in uncorrelated states, the phase uncertainty is limited by the standard quantum limit (SQL) $\Delta\phi^{2}\geq\Delta\phi_{\mathrm{SQL}}^{2}=1/N$ [2], such that $\Delta\omega(T)^{2}\geq(\Delta\omega_{\mathrm{SQL}}(T))^{2}=1/\tau TN$. This bound can be saturated with atoms prepared in a coherent spin state (CSS) and measurement of the spin projection along a suitable direction. A dramatic improvement can, in principle, be gained by preparing atoms in a GHZ state $|\mathrm{GHZ}\rangle=\left(|\negmedspace\uparrow\rangle^{\otimes N}+|\negmedspace\downarrow\rangle^{\otimes N}\right)/\sqrt{2}$ and measuring the parity $\Pi=(-1)^{N}\sigma_{x}^{\otimes N}$ after phase imprint during a Ramsey interrogation, as initially proposed by Wineland et al. [72]. This strategy, to which we refer in the following as ‘parity-GHZ’ protocol, surpasses the SQL and saturates the Heisenberg limit (HL) $(\Delta\phi_{\mathrm{HL}})^{2}=1/N^{2}$ [2]. The improvement is due to the signal in Ramsey interference oscillating $N$ times faster, thus enhancing the sensitivity by exactly this factor.
@@
For a given probe state, the quantum Cramér-Rao bound (QCRB) represents a lower limit in phase estimation, i.e. depicts the minimal phase uncertainty with respect to all conceivable measurement and estimation schemes [2]. Indeed, performing a parity measurement on the GHZ state in the presence of spontaneous decay is not optimal, i.e. does not saturate the QCRB. We show in the Supplementary Materials that the QCRB for the phase uncertainty is given by
@@
The corresponding bound for the optimal frequency uncertainty follows from minimizing with respect to the interrogation time, $(\Delta\omega_{\mathrm{QCRB}}^{\mathrm{GHZ}})^{2}=\min_{T}(\Delta\phi_{\mathrm{QCRB}}^{\mathrm{GHZ}}(T))^{2}/T\tau$. For vanishing spontaneous emission, $\Gamma=0$, this reproduces the result of Huelga et al. [43] that the QCRB equals the SQL. However, for $\Gamma>0$ one finds $(\Delta\omega_{\mathrm{QCRB}}^{\mathrm{GHZ}})^{2}<(\Delta\omega_{\mathrm{SQL}})^{2}$ which is illustrated in Fig. 2A. Somewhat surprisingly, GHZ states do admit gains beyond the SQL when the relevant decoherence process is spontaneous emission rather than dephasing noise. Since these gains are not realized by parity measurements, the question arises as to which alternative measurements do saturate the QCRB of the GHZ state (3).
@@
$\mathcal{U}_{\mathrm{GHZ}}=\begin{cases}\mathcal{T}_{x}(\pi)&\text{if }N\text{ is even}\\ \mathcal{R}_{x}\left(\frac{\pi}{2}\right)\mathcal{T}_{x}(\pi)&\text{if }N\text{ is odd}\end{cases}$ (4)
where $\mathcal{T}_{x}(\mu)=\exp\left(-i\frac{\mu}{2}S_{x}^{2}\right)$ denotes an OAT operation with twisting strength $\mu=\pi$ along $x$, whereas $\mathcal{R}_{x}(\theta)$ represents a rotation around the $x$-axis by an angle $\theta=\frac{\pi}{2}$. We note that the unitary $\mathcal{U}_{\mathrm{GHZ}}$ also corresponds to the operation that may be used to generate the GHZ state from the ground state $|\negmedspace\!\downarrow\rangle^{\otimes N}$ initially. Subsequently, atoms are measured projectively along $z$, effectively measuring the operator $X=\mathcal{U}_{\mathrm{GHZ}}\,S_{z}\,\mathcal{U}_{\mathrm{GHZ}}^{\dagger}$ with outcomes $x\in\{-\frac{N}{2},\ldots,\frac{N}{2}\}$. This procedure corresponds to the one explored in the experiment reported by Leibfried et al. [66] and essentially implements an exact Loschmidt echo [73], since the state preparation with $\mathcal{U}_{\mathrm{GHZ}}$ is complemented by the corresponding adjoint transformation $\mathcal{U}_{\mathrm{GHZ}}^{\dagger}$ for the measurement. Any such measurement has to be accompanied by a suitable rule for estimating the phase $\phi$ for a given outcome $x$. We can show that the nonlinear estimator
$\phi_{\mathrm{est}}(x,T)=\begin{cases}\pm\mathrm{e}^{\frac{(\Gamma+\gamma)}{2}NT}&\text{if }x=\pm\frac{N}{2}\\ 0&\text{else}\end{cases}$ (5)
@@
The fact that this measurement and estimation strategy performs well under spontaneous emission can be understood from the two following reasons: Firstly, measurement outcomes $x\neq\pm\frac{N}{2}$ can occur only if at least one spontaneous emission event happened, such that selecting the $x=\pm\frac{N}{2}$ events, as is done by the estimator (5), in essence filters out all such cases. Secondly, and quite remarkably, for the particular measurement $X$ introduced above, the conditional probabilities $P(x|\phi,T)$ turn out to be independent of $\phi$ for all $x\neq\pm\frac{N}{2}$ (cf. Methods). Thus, these cases don’t deliver any information on $\phi$. Together, these two features allow to implement an error detection and mitigation scheme [74] tailored to frequency metrology. On the one hand, outcomes $x\neq\pm\frac{N}{2}$ can be considered as heralded errors signifying an unsuccessful, decohered Ramsey interrogation. Conversely, outcomes $x=\pm\frac{N}{2}$ signal a no-jump dynamics (cf. Supplementary Materials) delivering maximal phase information. On the other hand, exclusively selecting these events during (classical) post-processing results in enhanced sensitivity. We therefore refer to this scheme as ‘heralded-GHZ’ protocol in the following.
@@
Figure 2: Enhancement of the optimal GHZ protocol compared to the SQL. (A) Generic frequency estimation uncertainty $\Delta\omega(T)$ in the presence of spontaneous decay for $N=8$ scaled to the SQL (2), $\gamma=0$. Symbols indicate the optimal interrogation time $T_{\mathrm{min}}$. (B) Relative frequency estimation uncertainty at the optimal interrogation time $T_{\mathrm{min}}$ for the indicated protocol types (see main text). The ratio $\Delta\omega/\Delta\omega_{\mathrm{SQL}}$ is independent of the decay rate $\Gamma$ due to the comparison with the SQL (triangle in (a)). Square and diamond symbols correspond to the minima in (a). The inset illustrates the asymptotic scaling for large ensembles $N$, using the same axes and protocols.
@@
Optimization over all entangled initial states, measurements, and estimators determines the performance of an optimal quantum interferometer (OQI). In the absence of decoherence, the OQI achieves the Heisenberg limit of sensitivity. Unfortunately, in the presence of decoherence, no general expressions for the OQI sensitivity for arbitrary ensemble sizes are available, but they rather require complex optimization procedures. An algorithm presented in Ref. [58] iteratively optimizes the initial probe state and the observable, which can be done efficiently at least for up to $N=30$ atoms. However, numerical optimization becomes challenging with increasing ensemble size $N$ (cf. Supplementary Materials). In the asymptotic limit ($N\gg 1$), lower bounds can be derived for several noise processes as discussed in Refs. [42, 41, 43, 75, 76, 77]. The ultimate lower bound in the presence of dephasing [43, 42, 41] and spontaneous decay [77] is given by
@@
which yields a maximal improvement of $1/\mathrm{e}$ over the SQL (2), corresponding to a gain of 4.3 dB, and is asymptotically saturated by SSS (cf. Supplementary Materials). As a consequence, we will focus on small ensembles where the OQI can be evaluated numerically and benchmark the heralded-GHZ protocol against SSS generated by OAT since they are optimal in the asymptotic limit. In the following, we will consider setups where spontaneous decay is the main limiting decoherence process, and neglect dephasing noise assuming $\gamma=0$.
@@
In Fig. 2B, we present the sensitivity of frequency estimation for a given particle number $N$, optimized in each case with respect to the interrogation time $T$ and scaled to the SQL. The conclusions we draw from Fig. 2B are: (i) The heralded-GHZ protocol shows a substantial enhancement over the SQL for all $N$. In particular, a constant gain of 1.8 dB with regard to the the SQL is achieved for $N\geq 4$ and thus, no loss in improvement is observed for large $N$. Especially considering the scaling of small and intermediate ensembles, the resulting increase in sensitivity is quite remarkable. (ii) The importance of the nonlinear estimator (5) becomes evident since otherwise (denoted as ‘linear-GHZ’ protocol), i.e. using the standard linear estimator, the maximal gain is obtained for $N=3$, whereas for larger ensembles the sensitivity converges to the SQL (cf. Methods). (iii) Although the heralded-GHZ protocol provides a substantial enhancement compared to the SQL, it does not saturate the OQI, which likewise represents a constant improvement with respect to the SQL in the regime ($N\leq 30$) where the numerical evaluation is feasible (cf. Supplementary Materials). Furthermore, in this regime the heralded-GHZ protocol achieves sensitivities relatively close to the OQI, despite its low complexity. (iv) For larger ensembles ($N\geq 42$), SSS achieve a higher sensitivity than the heralded-GHZ protocol and ultimately approximate the lower bound (6) asymptotically. Consequently, the gain of the OQI over the SQL will increase likewise for larger ensembles. In summary, this implies that GHZ states can be used to overcome the SQL in the presence of spontaneous emission and in addition perform close to the OQI for small and moderate ensemble sizes.
Saturating the OQI—
The gap between the heralded-GHZ protocol and the OQI (cf. Fig. 2B) raises the question of which protocol could be used to close it and what resources would be required to do so. The fact that the gap is independent of the number of particles suggests that this could be possible with a fixed protocol that varies little or not at all with $N$. Indeed, we have identified a particular interrogation scheme that reaches the level of the OQI for up to about 80 atoms. Surprisingly, no deep circuit depths are necessary, but a simple extension of the heralded-GHZ protocol involving one more twisting operation for state preparation is sufficient. The initial state
$|\psi_{\mathrm{in}}\rangle$ $=\mathcal{U}_{\mathrm{GHZ}}\,\mathcal{R}_{z}(\theta)\,\mathcal{U}_{\mathrm{GHZ}}|\negmedspace\downarrow\rangle^{\otimes N}=\alpha_{\theta}|\negmedspace\downarrow\rangle^{\otimes N}+\beta_{\theta}|\negmedspace\uparrow\rangle^{\otimes N}$ (7)
is generated by an additional OAT interaction $\mathcal{U}_{\mathrm{GHZ}}$ and rotation $\mathcal{R}_{z}(\theta)$ with optimal rotation angle $\theta$ depending on the ensemble size $N$ and the dimensionless parameter $\Gamma T$ (cf. Supplementary Materials). Essentially, the additional transformation only modifies the coefficients $\alpha_{\theta}$ and $\beta_{\theta}$ generating an unbalanced GHZ-like state (referred to as uGHZ in the following). In particular, the optimal rotation angle gives higher weight to the excited state, compensating for decay during the free evolution time. Due to its GHZ-like nature, the same measurement (4) and estimation scheme (5) are optimal as for the heralded-GHZ protocol and thus result in similar properties. We show in the Supplementary Materials that the corresponding phase estimation uncertainty reads
$(\Delta\phi(T))^{2}=\frac{e^{(\Gamma+\gamma)NT}}{4N^{2}}\left[1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}\right]^{2}.$ (8)
Fig. 2B shows that this ‘heralded-uGHZ’ protocol saturates the OQI for small ensemble sizes ($N\leq 4$) and achieves a constant gain of 2.25 dB compared to the SQL for $N\geq 6$, which is close to the OQI for intermediate $N$. Again, the asymptotically optimal SSS are advantageous for larger ensembles ($N>87$).
@@
and thus corresponds to the frequency estimation uncertainty rescaled with $\omega_{0}$. Consequently, we present our results in two ways: Lower $x$-axis and left $y$-axis refer to general frequency estimation where the uncertainty is rescaled to be independent of the particular averaging time $\tau$ and lifetime $t_{\mathrm{spont}}=1/\Gamma$. Thus, it allows for an application to several experimental setups and atomic species as long as spontaneous decay remains the dominating decoherence effect. Upper $x$-axis and right $y$-axis illustrate results in an atomic clock for the particular example of Ca+-ions with lifetime $t_{\mathrm{spont}}=\frac{1}{\Gamma}\simeq 1.1\,\mathrm{s}$ [79] and transition frequency $\omega_{0}=2\pi\nu_{0}\simeq 411.042\,\mathrm{THz}$ [15]. Furthermore, we consider frequency fluctuations corresponding to a state of the art clock laser. In particular, we assume a laser limited by flicker noise with coherence time $Z\simeq 7.5~{}\mathrm{s}\gg t_{\mathrm{spont}}$ [57]. Results of numerical simulations (symbols) in comparison to theoretical predictions (lines) of the investigated protocols are shown in Fig. 3 for $N=4$ (top) and $N=16$ (bottom) particles. As Monte Carlo simulations are stochastic processes, resulting stabilities fluctuate around the average value. Overall, numerical simulations of all discussed interrogation schemes show very good agreement with theoretical predictions. Therefore, all schemes, including the heralded-GHZ and heralded-uGHZ protocols in particular, are robust and thus suited for realistic scenarios as in the context of atomic clocks. At long interrogation times $T\gtrsim t_{\mathrm{spont}}$, fringe hops occur where the feedback loop passes to an adjacent fringe resulting in the clock running systematically wrong and consequently spoiling the clock stability. Moreover, the comparison of $N=4$ and $N=16$ already indicates the transition between the optimality of GHZ-like protocols and squeezing protocols.
Figure 3: Monte-Carlo simulations of the full feedback loop in an atomic clock. Simulation results are displayed by symbols while theoretical predictions are shown as lines. Data points denote the average over $10$ independent clock runs. Error bars indicate the corresponding standard deviations. In each clock run, $10^{7}$ cycles were performed and the corresponding Allan deviation evaluated. The presented values are obtained by extrapolating the Allan deviation at $\tau=1~{}\mathrm{s}$ based on the long term stability (9) at $\tau\gg 1$.
Conclusion & Outlook—